# Lifelong Graph Summarization with Neural Networks: 2012, 2022, and a Time Warp

### v1

@@ 6. Discussion
Our experiments show that the problem of graph summarization by vertex classification is hard when applied to a temporal graph with changing and unseen classes. The results from the snapshots in $2012$ and $2022$ (Figures 3 and 4) show that the classification of EQCs for $M_{AC1}$ is easier than for $M_{AC2}$. This is caused by the increased number of EQCs when more hops are considered in the summary model. Figures 2(a) and 2(d) shows that there are more unique EQCs for $M_{AC2}$ than for $M_{AC1}$. The EQCs in $M_{AC2}$ also change more rapidly. For $2022$, the snapshots have more edges than the ones we use from $2012$. Overall, this reduces the performance. The more EQCs the network needs to learn and the more they change, the worse the network performs in general on all tasks.
@@
We observe that no network has a positive forward or backward transfer between different snapshots. That means the networks perform better if they are trained on the task before being tested on it. The reason for the negative forward transfer is that new classes appear, and the reason for the negative backward transfer is that the neural network forgets classes. The results (Figures 3 and  4) confirm this statement, since a network always achieves its best accuracy on the snapshot it has been trained on. This observation is confirmed by $\Omega_{\mathrm{new}}$, the average of the accuracies on tasks after a network has been trained, which was higher than the other accuracies. One explanation is that the tasks the networks faced were very different, meaning that the EQCs of the snapshots change a lot, as can be seen in Figures 2(b), 2(c), 2(e), and 2(f).
The $\Omega_{\mathrm{base}}$ values in 2012 show that for this set of snapshots, the networks perform well on the first snapshot at each time. This is not the case for the networks for $M_{AC2}$ in 2022, where the values are below $0.5$. The networks do not perform as well as they could on the first snapshot. This shows that a network performs best on a task after being trained for it.
@@
The $ACC$ and the $\Omega_{\mathrm{all}}$ value of a network indicate how well a v performs on previous tasks. Figures 5 and 6 show how much a network forgets over time. A higher $ACC$ and $\Omega_{\mathrm{all}}$ value do not necessarily mean a network forgets less. A good example is the MLP ($1$-hop) of 2012, which has a better $ACC$ and $\Omega_{\mathrm{all}}$ value than the $2$-hop models, but the progress of forgetting is higher.
@@
We investigate whether the information provided from a $1$-hop neighborhood versus using the $2$-hop neighborhood is helpful. The measures show that the networks for $M_{AC2}$, independent of the used information, performed similarly (for both 2012 and 2022). Graph-MLP and GraphSAINT perform similarly to MLP for $M_{AC2}$, even though they have access to $2$-hop information for the classification task. While the $2$-hop information for GraphSAINT includes the path information between the vertices, the MLP can only use $1$-hop information. In summary, the results show that $2$-hop information does not improve the performance for a $2$-hop summary. While this seems to be counter-intuitive, an explanation is that for the $2$-hop summary, there are five to ten times more classes than in the $1$-hop summary since the number of paths grows exponentially with the number of hops (see Figures 2(a) and 2(d)). This results in many small EQCs, i. e., many classes with low support that are hard to train on (see Appendix B).
@@
Galke et al. (2023) differentiate between warm restart and cold restart in the context of lifelong learning. A warm restart is reusing network parameters from previous tasks, and retraining from scratch is called a cold restart. Our time warp experiment (Table 2) does not show a clear benefit of a warm restart compared to a cold restart over a ten-year range. In general, it shows that a warm restart does not hurt a network’s performance.
@@
Our networks may perform better if we trained them for more iterations. However, in Appendix A, we increased the number of training iterations by an order of magnitude and observed that it has only a small effect of less than $2$ points on the classification accuracy. If we continue training on the current task it may put the networks at more risk of forgetting the previous tasks. Furthermore, we use a hash function to compute the gold standard of the vertex labels, i. e., the EQCs. This could lead to a collision in the hash function. However, we chose the hash such that the probability is extremely low following Blasi et al. (Blasi et al., 2022).
7. Conclusion and Future Work
@@
We show that graph summarization for temporal graphs using a neural network is more challenging than the equivalent task on static graphs. The performance of the vertex classification in a temporal graph depends on the number of classes and the changes between timesteps. A problem with the summarization is that new EQCs appear and known ones disappear over time. We observe forgetting in lifelong graph summarization by vertex classification. We argue that the performance of a network should not only be measured with the accuracy of the current task but also other measures should be taken into account, like forward and backward transfer. The lifelong learning measures show that the networks for $M_{AC2}$ perform similarly. Our experiments also show that just reusing the network parameters does not necessarily improve the lifelong learning measures.
In future work, one may use different lifelong learning approaches to mitigate forgetting and increase backward and forward transfer for lifelong graph summarization. It may be interesting to investigate different summaries and the impact of the singleton EQCs, i. e., where only a single vertex is summarized, which we did not address here. It may also be interesting to measure the similarity of the predicted EQCs to the actual EQCs computed from a graph summary algorithm. In our setting, retraining is always the best option, but future work can also look into the question of when to retrain, as done by Galke et al. (Galke et al., 2023).
Acknowledgements.
@@
This research is co-funded by the CodeInspector project (No. 504226141) of the DFG, German Research Foundation. The authors acknowledge support by the state of Baden-Württemberg through bwHPC.
Appendix A Hyperparameter Optimization
The results of our hyperparameter optimization are shown in Table 3. Since the search is very expensive, we first perform a grid search with each network and configuration for $46$ iterations. Subsequently, we narrowed the search and chose the three best configurations of a network for each summary. We validate every neural network for $751$ more iterations, except MLP when using the graph summary model $M_{AC1}$. For MLP with $M_{AC1}$, we use only $121$ iterations because its batches contain more samples than for the graph neural networks. The results show that training for more iterations does not improve the performance.
Network Model LS LR D $\alpha$ $\tau$ Acc(1) Acc(2) Acc(3) Acc(4) Acc(5)
MLP $M_{AC1}$ 1,024 0.1 0.5 0.8481 0.8380 0.8357 0.8297 0.8440
MLP $M_{AC1}$ 1,024 0.01 0.5 0.7969 0.8398 0.8499 0.8496 0.8532
MLP $M_{AC1}$ 1,024 0.1 0.2 0.8378 0.8367 0.8293 0.8366 0.8367
MLP $M_{AC2}$ 1,024 0.01 0.5 0.5758 0.5775 0.5807 0.5832 0.5752
MLP $M_{AC2}$ 1,024 0.1 0.5 0.5479 0.5562 0.5524 0.5548 0.5458
MLP $M_{AC2}$ 1,024 0.1 0.2 0.5539 0.5633 0.5546 0.5543 0.5518
GraphSAINT $M_{AC2}$ 64 0.1 0.2 0.5654 0.5589 0.5714 0.5575 0.5621
GraphSAINT $M_{AC2}$ 32 0.1 0.2 0.5686 0.5639 0.5518 0.5559 0.5588
GraphSAINT $M_{AC2}$ 64 0.1 0 0.5736 0.5797 0.5720 0.5788 0.5686
GraphSAINT (Edges) $M_{AC2}$ 32 0.1 0 0.5558 0.5665 0.5550 0.5656 0.5540
GraphSAINT (Edges) $M_{AC2}$ 64 0.1 0 0.5409 0.5566 0.5481 0.5389 0.5352
GraphSAINT (Edges) $M_{AC2}$ 64 0.1 0.5 0.5350 0.5437 0.5223 0.5350 0.5342
Graph-MLP $M_{AC2}$ 64 0.01 0.2 1 2 0.6184 0.6250 0.6280 0.6180 0.6286
Graph-MLP $M_{AC2}$ 64 0.01 0 1 0.5 0.6222 0.6294 0.6253 0.6257 0.6269
Graph-MLP $M_{AC2}$ 64 0.01 0 10 0.5 0.6195 0.6228 0.6210 0.6246 0.6258
Table 3. Extensive search for parameters for the networks. Abbreviations: Layer Size (LS), Learning Rate (LR), and Dropout (D)). For MLP ($M_{AC1}$), the accuracy Acc($n$) is measured after the network has been trained for $1+n\cdot 24$ iterations and for the other networks after they have been trained for $1+n\cdot 150$ iterations. Best networks are in bold.
Appendix B Graph Summary Measures
To analyze the graphs and summaries more extensively, we exploit additional measures. We divide these measures into three categories. The first category is the unary measures. The second is binary measures, and the last is meta-measures.
B.1. Unary Measures
Unary measures only consider a summary $S_{t}$ at time $t$. The features of each vertex in the summary are the hash value that indicates its EQC, the degree, and the members of the EQC in the original graph. For each summary $S_{t}$, we have a function $ext_{t}()$ that returns for each vertex in the summary the number of vertices extensions in the original graph $G_{t}$. For the vertices, i.e., EQCs in the summary, $S_{t}$, we calculate the average size of EQCs as
(1) $\delta^{(1)}_{AvgSize}(S_{t})=\frac{\sum_{q\in V^{S}_{t}}ext_{t}(q)}{|V^{S}_{t}|},$
and the average number of edges/attributes per EQC as
(2) $\delta^{(1)}_{AvgE}(S_{t})=\frac{|E^{S}_{t}|}{|V^{S}_{t}|}.$
We also calculate distributions over the number of attributes of an EQC, over the number of members of an EQC, and over the number of usages of a predicate.
Figure 7 shows the average sizes of the EQCs as well as the average numbers of edges. Example distributions are depicted in Figures 8 and 9 for one snapshot in 2012 and 2022, respectively.
(a) Average number of attributes per EQC in 2012 ($1$-hop).
(b) Average number of members per EQC in 2012 ($1$-hop).
(c) Average number of attributes per EQC in 2012 ($2$-hop).
(d) Average number of members per EQC in 2012 ($1$-hop).
(e) Average number of attributes per EQC in 2022 ($1$-hop).
(f) Average number of members per EQC in 2022 ($1$-hop).
(g) Average number of attributes per EQC in 2022 ($2$-hop).
(h) Average number of members per EQC in 2022 ($2$-hop).
Figure 7. Average numbers for attributes (AvgE) and members (AvgSize) per EQC in the snapshots 2012 and 2022.
(a) Frequency of number of attributes per EQC (in descending order) for the $1$-hop attribute collection.
(b) Frequency of number of attributes per EQC (in descending order) for the $2$-hop attribute collection.
(c) Frequency of number of members of an EQCs (in descending order) for the $1$-hop attribute collection.
(d) Frequency of number of members of an EQCs (in descending order) for the $2$-hop attribute collection.
(e) Frequency how often a predicate is used in the original graph (in descending order).
Figure 8. Distributions measures on the summary of 2012-05-06
(a) Frequency of number of attributes per EQC (in descending order) for the $1$-hop attribute collection.
(b) Frequency of number of attributes per EQC (in descending order) for the $2$-hop attribute collection.
(c) Frequency of number of members of an EQCs (in descending order) for the $1$-hop attribute collection.
(d) Frequency of number of members of an EQCs (in descending order) for the $2$-hop attribute collection.
(e) Frequency how often a predicate is used in the original graph (in descending order).
Figure 9. Distributions measures on the summary of 2022-09-25
B.2. Binary Measures
Binary measures are used to compare a current summary $S_{t+1}$ to the previous summary $S_{t}$. It is used to measure the change that occurred between $S_{t+1}$ and $S_{t}$. The following measure is calculated with the Jaccard coefficient to calculate the change of vertices:
(3) $\delta^{(2)}_{JacV}(S_{t},S_{t+1})=1-\frac{|V^{S}_{t}\cap V^{S}_{t+1}|}{|V^{S}_{t}\cup V^{S}_{t+1}|}.$
We also calculate the Jensen–Shannon Divergence of $S_{t+1}$ and $S_{t}$. The Jensen–Shannon Divergence can be seen as the relative entropy between two probability distributions and is calculated as
(4) $\delta^{(2)}_{JS}(A,B)=D(A,B)+D(B,A)$
with
(5) $D(A,B)=\sum_{q\in V^{B}}P_{A}(q)\log_{2}\frac{P_{A}(q)}{P_{B}(q)}$
and $P_{A}(q)=\frac{ext_{t}(q)}{|V_{A}|},$ where $V_{A}$ are the vertices in A. If $ext_{t}(q)=0$, then $P_{A}(eq)=0$, and we consider the summand to be 0 as
$\lim_{x\rightarrow 0}x\log(x)=0.$
The changes in the binary measures are shown in Figure 10.
B.3. Meta-Measures
With meta-measures, we track how the summaries change over time, and in contrast to binary measures, it considers more than two summaries. During our experiments, we track whether new EQCs are appearing compared to the first snapshot and the previous one. We also track whether EQCs are reappearing or disappearing and how many EQCs have been seen until a certain snapshot. The meta-measures of our snapshots in 2012 and 2022 are depicted in Figure 11.
(a) Jaccard Coefficients in 2012.
(b) Jaccard Coefficients in 2022.
(c) Jensen-Shannon Divergences in 2012.
(d) Jensen-Shannon Divergences in 2022.
Figure 10. Jaccard Coefficients and Jensen-Shannon Divergences in the snapshots of 2012 and 2022.
(a) Changes of the EQCs compared to the first one in 2012 ($1$-hop).
(b) Changes of the EQCs compared to the previous one in 2012 ($1$-hop).
(c) Changes of the EQCs compared to the first one in 2012 ($2$-hop).
(d) Changes of the EQCs compared to the previous one in 2012 ($2$-hop).
(e) Changes of the EQCs compared to the first one in 2022 ($1$-hop).
(f) Changes of the EQCs compared to the previous one in 2022 ($1$-hop).
(g) Changes of the EQCs compared to the first one in 2022 ($2$-hop).
(h) Changes of the EQCs compared to the previous one in 2022 ($2$-hop).
(i) Number of EQCs in each snapshot of 2012.
(j) Number of all seen EQCs till a snapshot in 2012.
(k) Number of EQCs in each snapshot of 2022.
(l) Number of all seen EQCs till a snapshot in 2022.
Figure 11. Measuring the changes in the EQCs per snapshot in 2012 and 2022.
Appendix C Impact of Layer Normalization
We investigate the impact of normalizing a GCN, we normalize the GCN by using the degrees of the vertices. The adapted activation $h^{k}_{v}$ of each vertex $v$ in layer $k$ is calculated as follows:
(6) $h^{k}_{v}=\sigma\left(\sum_{u\in V_{s}}\frac{1}{\sqrt{\hat{d}_{v}\hat{d}_{u}}}\tilde{A}_{P}[v,u](W^{k-1})^{T}h_{u}^{(k-1)}\right)\,,$
where $\hat{d}_{v}$ is the degree of $v$ and $\hat{d}_{u}$ is the degree of $u$.
The results of the accuracy scores depicted in Figure 12 show that normalization does not make a big difference. This also is true of the forgetting scores, which are reported in Tables 4 and 5.
(a) GraphSAINT ($2$-hop)
(b) GraphSAINT ($2$-hop and normalized).
Figure 12. Accuracies comparison (Normalization vs. no normalization)
2012 GraphSAINT $2$-hop 0.376 -0.087 -0.043 0.824 0.445 0.672
2012 GraphSAINT $2$-hop and normalized 0.372 -0.091 -0.042 0.825 0.445 0.675
Table 4. Lifelong Learning Measures (Normalization vs. no normalization).
Year Network Type $F_{2}$ $F_{3}$ $F_{4}$ $F_{5}$ $F_{6}$ $F_{7}$ $F_{8}$ $F_{9}$ $F_{10}$
2012 GraphSAINT $2$-hop 0.010 0.042 0.071 0.032 0.037 0.080 0.044 0.071 0.087
2012 GraphSAINT $2$-hop and normalized 0.019 0.047 0.064 0.017 0.047 0.081 0.049 0.065 0.093
Table 5. Forgetting values (Normalization vs. no normalization).

### v2

@@ VI Discussion
Our experiments show that the problem of graph summarization by vertex classification is hard when applied to a temporal graph with changing and unseen classes. The results from the snapshots in $2012$ and $2022$ (Figures 3 and 4) show that the classification of EQCs for $M_{\mathrm{AC1}}$ is easier than for $M_{\mathrm{AC2}}$. This is caused by the increased number of EQCs when more hops are considered in the summary model. Figures 2(a) and 2(b) show that there are more unique EQCs for $M_{\mathrm{AC2}}$ than for $M_{\mathrm{AC1}}$. The EQCs in $M_{\mathrm{AC2}}$ also change more rapidly. The 2022 snapshots have more edges than the ones from 2012. Overall, this reduces the performance. The more EQCs the network needs to learn and the more they change, the worse the network performs in general on all tasks.
@@
We observe that no network has a positive forward or backward transfer between different snapshots. That means the networks perform better if they are trained on the task before being tested on it. The reason for the negative forward transfer is that new classes appear, and the reason for the negative backward transfer is that the neural network forgets classes. The results (Figures 3 and  4) confirm this statement, since a network that has been trained on $\mathcal{T}_{1},\dots,\mathcal{T}_{i}$ always achieves its best accuracy on $\mathcal{T}_{i}$. This observation is confirmed by $\Omega_{\mathrm{new}}$, the average of the accuracies on tasks after a network has been trained, which was higher than the other accuracies. One explanation is that the EQCs of the snapshots change a lot, as can be seen in Figures 2(c)2(f).
The $\Omega_{\mathrm{base}}$ values in 2012 show that for this set of snapshots, the networks perform well on the first snapshot at each time. This is not the case for the networks for $M_{\mathrm{AC2}}$ in 2022, where the values are below $0.5$. The networks do not perform as well as they did on the first snapshot. This shows that a network performs best on a task after being trained for it.
@@
The $ACC$ and the $\Omega_{\mathrm{all}}$ value of a network indicate how well a neural network performs on previous tasks. Figures 5 and 6 show how much a network forgets over time. A higher $ACC$ and $\Omega_{\mathrm{all}}$ value do not necessarily mean a network forgets less. A good example is the MLP ($1$-hop) of 2012, which has a better $ACC$ and $\Omega_{\mathrm{all}}$ value than the $2$-hop models, but the forgetting rate is higher.
@@
We investigate whether the information provided from a $1$-hop neighborhood versus using the $2$-hop neighborhood is helpful. The measures show that the networks for $M_{\mathrm{AC2}}$, independent of the information used, performed similarly, and this held in both 2012 and 2022. Graph-MLP and GCN perform similarly to MLP for $M_{\mathrm{AC2}}$, even though they have access to $2$-hop information for the classification task and MLP does not. In summary, the results show that $2$-hop information does not improve the performance for $M_{\mathrm{AC2}}$. While this seems to be counter-intuitive, an explanation is that for the $2$-hop summary, there are five to ten times more classes than in the $1$-hop summary since the number of paths grows exponentially with the number of hops (see Figures 2(a) and 2(b)). This results in many small EQCs, i. e., many classes with low support that are hard to train on.
@@
Galke et al. [21] differentiate between warm restart and cold restart in the context of lifelong learning. A warm restart is reusing network parameters from previous tasks, and retraining from scratch is called a cold restart. Our time warp experiment (Table II) does not show a clear benefit of a warm restart compared to a cold restart over a ten-year range. However, it shows that a warm restart does not hurt a network’s performance.
@@
The networks may perform better if we trained them for more iterations. However, when we increase the number of training iterations by an order of magnitude, we observe an improvement of only $2$ points on the classification accuracy. Training more on the current task may also increase the forgetting of previous tasks. Furthermore, we use a hash function to compute the gold standard of the vertex labels, i. e., the EQCs. There could be collisions in the hash function. However, we chose the hash such that the probability is very low following Blasi et al. [6].
VII Conclusion
@@
We show that graph summarization for temporal graphs using a neural network is more challenging than the equivalent task on static graphs. The performance of the vertex classification in a temporal graph depends on the number of classes and the changes between timesteps. A problem with the summarization is that many new EQCs appear, and known ones disappear over time. We observe forgetting in lifelong graph summarization by vertex classification. We also observe that the performance of a network should not only be measured by the accuracy of the current task but also by other measures, like forward and backward transfer, which should be taken into account. The measures show that the networks for $M_{\mathrm{AC2}}$ perform similarly. With the time warp experiments, we see that reusing the network parameters does not necessarily improve lifelong learning measures.
Acknowledgements
@@
This paper is the result of the first author’s master’s project in data science. This research is co-funded by the CodeInspector project (No. 504226141) of the DFG, German Research Foundation, and the 2LIKE project by the German Federal Ministry of Education and Research (BMBF) and the Ministry of Science, Research and the Arts Baden-Württemberg within the funding line Artificial Intelligence in Higher Education. The authors acknowledge support by the state of Baden-Württemberg through bwHPC.